OpenStudy (anonymous):

A recursive function is shown below: f(1) = 4 and f(n) = f(n - 1) - 8; n > 1 Which of the following lists the terms in the sequence defined by this recursive function? 4, -4, -12, -20, -28, ... 4, 4, 12, 20, 28, ... 4, 12, 20, 28, 36, ... 4, -12, -20, -28, -36, ...

OpenStudy (anonymous):

@jim_thompson5910 Can you help me again?

jimthompson5910 (jim_thompson5910):

are you familiar with the notation f(n) = f(n - 1) - 8 ?

OpenStudy (anonymous):

no

jimthompson5910 (jim_thompson5910):

f(n) is the nth term f(n-1) is the term just before the nth term

OpenStudy (anonymous):

I have my notes with me but they are no help

jimthompson5910 (jim_thompson5910):

saying f(n) = f(n - 1) - 8 means "nth term = (term before nth term) - 8"

jimthompson5910 (jim_thompson5910):

if you are given a term, you can find the next term by subtracting 8

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

so you would subtract 8 from 4?

jimthompson5910 (jim_thompson5910):

correct

OpenStudy (anonymous):

so it would be 4,-12,...

OpenStudy (anonymous):

so then you would keep going ?

jimthompson5910 (jim_thompson5910):

first term is 4 second term is 4-8 = -4 third term is -4-8 = -12 etc etc notice how I'm subtracting 8 from each term

OpenStudy (anonymous):

So it would be 4,-12,-20,-28,-36...?

jimthompson5910 (jim_thompson5910):

yes

OpenStudy (anonymous):

thank you :D

jimthompson5910 (jim_thompson5910):

yw